Please answer the following question completely and correctly in order to be rated. Thank you.

You want to express protein xin E. coli cells and designed an experiment where you litated the DNA sequence from protein x (a) into the pTrcHiSA expression vector (b) using the EcoRI and Hindlll restriction sites. You then transformed the ligation (containing both vector only and vector inser nto coli cells. After extract performed a ng the DNA from multiple single colonies, you the gel restriction digest with Pvul. Based on the restriction maps provided in (a & b), use provided in (c) to draw the bands you would expect to see for the vector only digest and for the vector containing the Protein X sequence pTrcHis- 412 Pvul Ladder Proteinx sequence from 1 to 686 8 pts). 0.5 650 Apal (1) 4414 bp 1499 Pvul (1)
Answer #1

we will get one band near around 4.5kb in vector alone digested with Pvu1.

when we digest the vector+sequence we will get two bands of 1kb and 4kb length.

digestion with Ecor1 and Hind3 will remove 107 bp from vector and ligation of sequence will add 606 bp in vector. so overall 4414+606-107bp= 4913 bp. (606 bp addition because when you digest sequence with EcoR1 and Hind3 it will cut at seq 72 bp and 678 bp. so 678-72=606 bp.

now Pvu site is 340 bp away from Ecor1. so total base= 552+340= 892 bp is the new cut site of Pvu1 in vector+ sequence. now we have 266 bp extra (340+266=606bp).

distance between Hind3 site and Pvu1 site in vector is 1499-659=840bp. but we have 266 bp remaining of sequence. so total distance between Pvu1 sites of vector and sequence= 840+266=1106 bp.

and 4913-1106bp= 3807bp which is approx 4kb.

so we'll get two band of approx 1kb and 4 kb.